Determine the x-coordinate of the centroid of the shaded area.

19 — — — — — у a І, b2 ІІ х
Answer #1

dA х X а

Equation of the bottom straight line can be written as -+-=1 a b On rearranging y=b(1-3) = (a – x) = fı(1) Rearranging the eq

b A= a -sin -1 .x²+ b -[ax): + che le 02) - Bla?] + ba a 2 a a 0 ba (Tt - 2) 11 Centroid of the given area can be calculated